OpenStudy (anonymous):

Can you please show me how to solve for y? 3/4x-1/6y=x-2

hero (hero):

What the....?

hero (hero):

Didn't we go over this already?

OpenStudy (anonymous):

Sorry, I lost it

hero (hero):

But you have it now, right?

OpenStudy (anonymous):

I was not able to sign back in last night

hero (hero):

You deleted the other one.

OpenStudy (anonymous):

I was not able to sign back in last night

OpenStudy (anonymous):

yes

hero (hero):

any reason why?

jhonyy9 (jhonyy9):

3/4x - 1/6y=x-2 9x-2y=12(x-2) 9x-2y=12x-24 -3x-2y=-24 24-3x=2y y=(24-3x)/2

hero (hero):

that's the same exact solution I came up with

OpenStudy (anonymous):

I was not able to sign back in last night

OpenStudy (anonymous):

Thanks to both of you =D

hero (hero):

jaramillotina, that's not fair that you deleted the other one. I can't understand why you did it

hero (hero):

No, you deleted the other one.

OpenStudy (anonymous):

I was not able to sign back in last night

OpenStudy (anonymous):

i wrote it down, but I don't have it with me.

hero (hero):

You shouldn't have deleted it

OpenStudy (anonymous):

I was not able to sign back in last night

OpenStudy (anonymous):

I know that now. I am new on this site. Do you have problems with the server not being available. I also have to reset my password everytime I sign back in.

hero (hero):

I still don't understand what motivated you to delete it from the beginning. There was no reason to.

OpenStudy (anonymous):

I was not able to sign back in last night

OpenStudy (anonymous):

I was deleting them as I got answers.

OpenStudy (anonymous):

I was not able to sign back in last night

OpenStudy (anonymous):

My lunch break is over, but thank you Hero for yesterday. Don't be upset.

OpenStudy (anonymous):

3/4x-1/6y=x-2 3/4x - 1/6y = x - 2 Subtract 3/4x from both sides. -1/6y = x - 2 - 3/4x -1/6y = 1/4x - 2 Divide through by -1 1/6y = -1/4x + 2 Multiply through by 6. y = -1/4x(6) + 2(6) y = -6/4x + 12 y = -3/2x + 12 Hope this helps.